# Unraveling the Interplay between Carryover Effects and Reward Autocorrelations in Switchback Experiments

### v6

@@ 3 Preliminaries
  • For a general $22\leq m\leq T/$ 2, the policy remains constant for $m$ time steps and then switches, which can be mathematically represented as $A_{mt-m+1}=A_{mt-m+2}=\ldots=A_{mt}=1-A_{mt+1}=\ldots=1-A_{mt+m}$;
@@
where ${\mathbb{E}}^{1}$ and ${\mathbb{E}}^{0}$ represent the expectation when the new policy (coded as $1$) and the baseline policy (coded as 00) are applied at all times, respectively. This metric is particularly relevant for A/B testing in RL (Tang et al., 2022).
@@ 4.1 Methods
where the product $\Phi_{t-1}\ldots\Phi_{k+1}$ is treated as an identity matrix if $t-1<k+1$. The first term on the right-hand side (RHS) of (2) represents the direct effect of actions on immediate rewards, while the latter term accounts for the delayed or carryover effects of previous actions.
@@
[…]
@@ 4.2 Theoretical analysis
Let $\sigma_{e}(t_{1},t_{2})=\sigma^{2}\rho^{|t_{1}-t_{2}|}$ for some $-1<\rho<1$ and $\sigma^{2}>0$. For sufficiently large $T$, the first term in (5) becomes asymptotically equivalently to
@@
Assume $\sigma_{e}(t_{1},t_{2})=\sigma^{2}[\rho\mathbb{I}(t_{1}\neq t_{2})+\mathbb{I}(t_{1}=t_{2})]$ for some $-1<\rho<1$ and $\sigma^{2}>0$. Then the first term in (5) equals
@@ Appendix A Implementation Details
From this, we compute $\widehat{V}_{t:t}^{a,m}(\tilde{s})$ as $\varphi^{\top}(\tilde{s})\widehat{\theta}_{0,a,m}$. Next, we sequentially compute $\{V_{t-1:t}^{a}\}_{t}$,$\{V_{t-2:t}^{a}\}_{t}$, and so forth. Specifically, for each $j=1,\ldots,T-1$, we recursively solve the following estimating equation,
@@
This leads to the derivation of $\widehat{\theta}_{j,a,m}$, based on which we set $\widehat{V}_{t-j:t}^{a,m}(\tilde{s})$ to $\varphi^{\top}(\tilde{s})\widehat{\theta}_{j,a,m}$. Finally, we set the ATE estimator to
@@ Appendix C Additional Experimental Results
$\begin{split}&\{\alpha_{t}\}_{t}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\begin{cases}U[-1,-0.5]&\text{ with probability 0.5 }\\ U[0.5,1]&\text{ with probability 0.5 }\end{cases},\{\beta_{t}^{(j)}\}_{t,j}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\begin{cases}U[-0.3,-0.1]&\text{ with probability 0.5 }\\ U[0.1,0.3]&\text{ with probability 0.5 }\end{cases},\\ &\{\phi_{t}^{(j)}\}_{t,j}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\begin{cases}U[-1,-0.5]&\text{ with probability 0.5 }\\ U[0.5,1]&\text{ with probability 0.5 }\end{cases}.\end{split}$
@@ D.2 Theorem 1
  • For the OLS estimator, under Assumptions 15, its reminder term $=O(n^{-3/2}R_{\max}^{2}\log(nT))$.
@@ D.3 Proof of Theorem 1-OLS and Corollaries 13
and vec()($\bullet$)( ∙ ) denotes the operator that vectorize a matrix into a row vector.
@@
Consider the leading term first. Recall that (33) obtains a closed-form expression for $\Sigma_{t}^{*-1}$. With some calculations, it can be shown that the last row of $\Sigma_{t}^{*-1}$ equals $(-2,0,4)$. This leads to
the last entry of the leading $\textrm{the last entry of the leading term}=\frac{4}{n}\sum_{i=1}^{n}(A_{i,t}-\frac{1}{2})e_{i,t}.$ (45)
@@
the reminder $\textrm{the reminder term}=O\Big{(}\|\frac{\delta}{n}\sum_{i=1}^{n}Z_{i,t}e_{i,t}\|_{2}\Big{)}+O\Big{(}\frac{R_{\max}\log(nT)}{n}\Big{)},$
@@
To ease notation, let $c_{t}$ denote the $d$-dimensional vector $[\sum_{k=t}^{T-1}\beta_{k+1}^{\top}(\Phi_{k}\cdots\Phi_{t+1})]^{\top}$. According to the Lemma 2, with a sufficiently large $\kappa>0$, the asymptotic MSE of the ATE estimator is given by
@@
$&\textrm{MSE}(\textrm{ATE}_{\text{SB}}^{(m)})=\frac{16}{nT^{2}}{\mbox{Var}}\Big{[}\sum_{t=1}^{T}{(A_{t}-\frac{1}{2})e_{t}}\Big{]}+\frac{16}{nT^{2}}{\mbox{Var}}\Big{[}\sum_{t=1}^{T-1}{(A_{t}-\frac{1}{2})c_{t}^{\top}E_{t}}\Big{]}\\ +&\mathbb{E}(I+I^{*})^{2}+2\mathbb{E}(I+I^{*})\Big{[}\frac{4}{nT}\sum_{i=1}^{n}\sum_{t=1}^{T}{(A_{i,t}-\frac{1}{2})e_{i,t}}\Big{]}+2\mathbb{E}(I+I^{*})\Big{[}\frac{4}{nT}\sum_{i=1}^{n}\sum_{t=1}^{T-1}{(A_{i,t}-\frac{1}{2})}c_{t}^{\top}E_{i,t}\Big{]}.\end{aligned}$ (57)
@@
Finally, we analyze the first term on the RHS of the first line of (57). Consider a given $m$-switchback design. For any integers $1\leq t_{1}\leq t_{2}\leq T$, we represent them as $t_{1}=l_{1}+k_{1}m$ and $t_{2}=l_{2}+k_{2}m$ such that $1\leq l_{1},l_{2}\leq m$ and $0\leq k_{1},k_{2}<T/m$. By definition, ${\mbox{Cov}}(A_{t_{1}},A_{t_{2}})$ equals $0.25$ if $k_{2}-k_{1}$ is even and $-0.25$ otherwise. Then the first term on the RHS of (57) can thus be represented as
@@
Recall that under the autoregressive covariance structure, $\sigma_{e}(t_{1},t_{2})=\sigma^{2}\rho^{t_{2}-t_{1}}$ for some $10<\rho<$ 1 and any $t_{1}\leq t_{2}$. It follows that
@@
where $[z]_{+}=\max(z,0)$ for any $z\in\mathbb{R}$. Accordingly, we have
@@ D.4 Proofs of Theorem 1-DRL and LSTD
$\frac{1}{T^{2}}\sum_{a=0}^{1}\sum_{t=1}^{T}{\mbox{Var}}\Big{[}\omega_{t}^{a,m}(A_{t},S_{t})\epsilon_{t}^{a}\Big{]}&=\frac{4}{T^{2}}\sum_{a=0}^{1}\sum_{t=1}^{T}{\mathbb{E}}\frac{[p_{t}^{a}(S_{t})]^{2}\mathbb{I}(A_{t}=a)\sigma^{2}_{\epsilon,t}(a,S_{t})}{[p_{t}^{m}(S_{t}|a)]^{2}}\\ &=\frac{2}{T^{2}}\sum_{a=0}^{1}\sum_{t=1}^{T}\sum_{s\in\mathcal{S}}\frac{[p_{t}^{a}(s)]^{2}\sigma^{2}_{\epsilon,t}(a,s)}{p_{t}^{m}(s|a)},\end{aligned}$ (65)
@@
×$\times[\widehat{V}_{t+1,-k}^{a,m}(S_{i,t+1})-\widehat{V}_{t,-k}^{a,m}(S_{i,t})-V_{t+1}^{a}(S_{i,t+1})+V_{t}^{a}(S_{i,t})].$
@@
×$\times[\widehat{V}_{t+1}^{a,m}(S_{i,t+1})-\widehat{V}_{t}^{a,m}(S_{i,t})-V_{t+1}^{a}(S_{i,t+1})+V_{t}^{a}(S_{i,t})].$
@@
where $\omega_{t}^{a,m*}(s)=\varphi_{t}^{\top}(s)\alpha_{t,a,m}^{*}$, and $\alpha_{t,a,m}^{*}:=\min_{\alpha_{t,a,m}\in\mathbb{R}^{L}}{\mathbb{E}}(\omega_{t}^{a,m}(S_{t})-\varphi_{t}^{\top}(S_{t})\alpha_{t,a,m})^{2}$ Then (74) is equal to
@@
$J_{4b}$ $:=\frac{2}{nT}\sum_{i=1}^{n}\sum_{t=1}^{T}\sum_{a=0}^{1}(-1)^{a+1}\mathbb{I}(A_{i,t}=a)\left[\widehat{\omega}_{t}^{a,m}(S_{i,t})-\omega_{t}^{a,m*}(S_{i,t})\right]\varepsilon_{i,t}^{a}.$
@@
[…]
@@
[…]

### v7

@@ 3 Preliminaries
  • For a general $2\leq m\leq T/2$, the policy remains constant for $m$ time steps and then switches, which can be mathematically represented as $A_{mt-m+1}=A_{mt-m+2}=\ldots=A_{mt}=1-A_{mt+1}=\ldots=1-A_{mt+m}$;
@@
where ${\mathbb{E}}^{1}$ and ${\mathbb{E}}^{0}$ represent the expectation when the new policy (coded as $1$) and the baseline policy (coded as 0) are applied at all times, respectively. This metric is particularly relevant for A/B testing in RL (Tang et al., 2022).
@@ 4.1 Methods
where the product $\Phi_{t-1}\ldots\Phi_{k+1}$ is treated as an identity matrix if t1<k+1t-1<k+1. The first term on the right-hand side (RHS) of (2) represents the direct effect of actions on immediate rewards, while the latter term accounts for the delayed or carryover effects of previous actions.
@@
$\begin{aligned} &\frac{1}{n}\sum_{i=1}^{n}\varphi_{t}(S_{i,t})\mathbb{I}(A_{i,t}=a)\Big{[}R_{i,t}\\ &+\varphi_{t+1}^{\top}(S_{i,t+1})\widehat{\theta}_{t+1,a,m}-\varphi_{t}^{\top}(S_{i,t})\widehat{\theta}_{t,a,m}\Big{]}=0.\end{aligned}$ (4)
@@ 4.2 Theoretical analysis
Let $\sigma_{e}(t_{1},t_{2})=\sigma^{2}\rho^{|t_{1}-t_{2}|}$ for some $1<\rho<1$ and $\sigma^{2}>0$. For sufficiently large $T$, the first term in (5) becomes asymptotically equivalently to
@@
Assume $\sigma_{e}(t_{1},t_{2})=\sigma^{2}[\rho\mathbb{I}(t_{1}\neq t_{2})+\mathbb{I}(t_{1}=t_{2})]$ for some $1<\rho<1$ and $\sigma^{2}>0$. Then the first term in (5) equals
@@ Appendix A Implementation Details
[…]
@@
[…]
@@ Appendix C Additional Experimental Results
[…]
@@ D.2 Theorem 1
  • For the OLS estimator, under Assumptions 15, its reminder term $O(n^{-3/2}R_{\max}^{2}\log(nT))$.
@@ D.3 Proof of Theorem 1-OLS and Corollaries 13
and vec()($\bullet$) denotes the operator that vectorize a matrix into a row vector.
@@
Consider the leading term first. Recall that (33) obtains a closed-form expression for $\Sigma_{t}^{*-1}$. With some calculations, it can be shown that the last row of $\Sigma_{t}^{*-1}$ equals (2,0,4)(-2,0,4). This leads to
$\textrm{the last entry of the leading term}=\frac{4}{n}\sum_{i=1}^{n}(A_{i,t}-\frac{1}{2})e_{i,t}.$ (45)
@@
$\textrm{the reminder term}=O\Big{(}\|\frac{\delta}{n}\sum_{i=1}^{n}Z_{i,t}e_{i,t}\|_{2}\Big{)}+O\Big{(}\frac{R_{\max}\log(nT)}{n}\Big{)},$
@@
To ease notation, let $c_{t}$ denote the $d$-dimensional vector $\sum_{k=t}^{T-1}\beta_{k+1}^{\top}(\Phi_{k}\cdots\Phi_{t+1})]^{\top}$. According to the Lemma 2, with a sufficiently large $\kappa>0$, the asymptotic MSE of the ATE estimator is given by
@@
$\begin{aligned} &\textrm{MSE}(\textrm{ATE}_{\text{SB}}^{(m)})=\frac{16}{nT^{2}}{\mbox{Var}}\Big{[}\sum_{t=1}^{T}{(A_{t}-\frac{1}{2})e_{t}}\Big{]}+\frac{16}{nT^{2}}{\mbox{Var}}\Big{[}\sum_{t=1}^{T-1}{(A_{t}-\frac{1}{2})c_{t}^{\top}E_{t}}\Big{]}\\ +&\mathbb{E}(I+I^{*})^{2}+2\mathbb{E}(I+I^{*})\Big{[}\frac{4}{nT}\sum_{i=1}^{n}\sum_{t=1}^{T}{(A_{i,t}-\frac{1}{2})e_{i,t}}\Big{]}+2\mathbb{E}(I+I^{*})\Big{[}\frac{4}{nT}\sum_{i=1}^{n}\sum_{t=1}^{T-1}{(A_{i,t}-\frac{1}{2})}c_{t}^{\top}E_{i,t}\Big{]}.\end{aligned}$ (57)
@@
Finally, we analyze the first term on the RHS of the first line of (57). Consider a given $m$-switchback design. For any integers $1\leq t_{1}\leq t_{2}\leq T$, we represent them as $t_{1}=l_{1}+k_{1}m$ and $t_{2}=l_{2}+k_{2}m$ such that $1\leq l_{1},l_{2}\leq m$ and $0\leq k_{1},k_{2}<T/m$. By definition, ${\mbox{Cov}}(A_{t_{1}},A_{t_{2}})$ equals $0.25$ if $k_{2}-k_{1}$ is even and 0.25-0.25 otherwise. Then the first term on the RHS of (57) can thus be represented as
@@
Recall that under the autoregressive covariance structure, $\sigma_{e}(t_{1},t_{2})=\sigma^{2}\rho^{t_{2}-t_{1}}$ for some $0<\rho<1$ and any $t_{1}\leq t_{2}$. It follows that
@@
where $z]_{+}=\max(z,0)$ for any $z\in\mathbb{R}$. Accordingly, we have
@@ D.4 Proofs of Theorem 1-DRL and LSTD
$\begin{aligned} \frac{1}{T^{2}}\sum_{a=0}^{1}\sum_{t=1}^{T}{\mbox{Var}}\Big{[}\omega_{t}^{a,m}(A_{t},S_{t})\epsilon_{t}^{a}\Big{]}&=\frac{4}{T^{2}}\sum_{a=0}^{1}\sum_{t=1}^{T}{\mathbb{E}}\frac{[p_{t}^{a}(S_{t})]^{2}\mathbb{I}(A_{t}=a)\sigma^{2}_{\epsilon,t}(a,S_{t})}{[p_{t}^{m}(S_{t}|a)]^{2}}\\ &=\frac{2}{T^{2}}\sum_{a=0}^{1}\sum_{t=1}^{T}\sum_{s\in\mathcal{S}}\frac{[p_{t}^{a}(s)]^{2}\sigma^{2}_{\epsilon,t}(a,s)}{p_{t}^{m}(s|a)},\end{aligned}$ (65)
@@
$\times[\widehat{V}_{t+1,-k}^{a,m}(S_{i,t+1})-\widehat{V}_{t,-k}^{a,m}(S_{i,t})-V_{t+1}^{a}(S_{i,t+1})+V_{t}^{a}(S_{i,t})].$
@@
$\times[\widehat{V}_{t+1}^{a,m}(S_{i,t+1})-\widehat{V}_{t}^{a,m}(S_{i,t})-V_{t+1}^{a}(S_{i,t+1})+V_{t}^{a}(S_{i,t})].$
@@
where $\omega_{t}^{a,m*}(s)=\varphi_{t}^{\top}(s)\alpha_{t,a,m}^{*}$, and $\alpha_{t,a,m}^{*}:=\min_{\alpha_{t,a,m}\in\mathbb{R}^{L}}{\mathbb{E}}(\omega_{t}^{a,m}(S_{t})-\varphi_{t}^{\top}(S_{t})\alpha_{t,a,m})^{2}$ . Then (74) is equal to
@@
$J_{4b}$ $=\frac{2}{nT}\sum_{i=1}^{n}\sum_{t=1}^{T}\sum_{a=0}^{1}(-1)^{a+1}\mathbb{I}(A_{i,t}=a)\left[\widehat{\omega}_{t}^{a,m}(S_{i,t})-\omega_{t}^{a,m*}(S_{i,t})\right]\varepsilon_{i,t}^{a}.$
@@
where the second inequality follows from the Cauchy-Schwarz inequality and that $\max_{t,a}|\varepsilon_{t}^{a}|=O(R_{\max})$. This leads to ${\mathbb{E}}(J_{4a}^{2})=O(n^{-1}R_{\max}^{2}\textrm{app\_err}_{\omega}^{2})$, where we denote $\textrm{app\_err}_{\omega}^{2}$ by $\max_{a,m,t}{\mathbb{E}}|\omega_{t}^{a,m}(S_{t})-\varphi_{t}^{\top}(S_{t})\alpha_{t,a,m}^{*}|^{2}$.
@@
where we denote $\textrm{est\_err}_{\omega}^{2}={\max_{a,t,m}{\mathbb{E}}\|\widehat{\alpha}_{t,a,m}-\alpha_{t,a,m}^{*}\|_{2}^{2}}$.